The father is 24 years older than his daughter.In 4 years, he will be thrice as old as his daughter.Find their present age.
Answer by lwsshak3(11628)

The father is 24 years older than his daughter.In 4 years, he will be thrice as old as his daughter.Find their present age.
***
x=present age of daughter
x+24=present age of father
..
in 4 yrs
x+4=age of daughter
x+24+4=x+28=age of father
x+28=3(x+4)
x+28=3x+12
2x=16
x=8
x+24=32
present age of daughter=8
present age of father=32